OpenStudy (anonymous):

What's 45% off of 230.45

OpenStudy (mertsj):

Are you aware that in mathematics, "of" means times?

OpenStudy (anonymous):

126.7475

OpenStudy (anonymous):

.45*230.45=103.7025- The amount that is minused from 230.45

OpenStudy (anonymous):

1. multiply first 45% by 230.45 so, 0.45 * 230.45 = 103.7025 2. since 45% of 230.45 is 103.7025, subtract this from the original number so, 230.45 - 103.7025 = 126.7475

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

Parth (parthkohli):

\(45\%\) off of \(230.45\) is just \(55\%\) of \(230.45\). As @Mertsj said, "of" means "times", and "\(55\%\)" means \(\frac{55}{100}\), so you need to find\[230 \times \dfrac{55}{100}\]